Grade 12th passAlgebra

Simplify:
((25) 2 × 73 )/ (83 × 7)
Simplify:
((25) 2 × 73 )/ (83 × 7)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

83 can be written as = (2 × 2 × 2)3
= (23)3
We have,
= ((25)2 × 73)/ ((23)3 × 7)
= (25 x 2 × 73)/ ((23 x 3 × 7) ... [∵(am)n = amn]
= (210 × 73)/ (29 × 7)
= (210 – 9× 73 – 1) ... [∵am ÷ an = am – n]
= 2 × 72
= 2 × 7 × 7
= 98

​Thanks